OpenStudy (fanduekisses):

What is the lim h->0 cos(3pi/2+h)-cos(3pi/2)/h

OpenStudy (turingtest):

I assume you have to solve this using nothing but the limits?

OpenStudy (misty1212):

oh i hope not!

OpenStudy (misty1212):

you have been asking calculus questions, implicit diff and so on much harder than this

OpenStudy (turingtest):

if you have to solve this with nothing but limits, use\[\cos(a+b)=\cos a\cos b-\sin a\sin b\]first

OpenStudy (turingtest):

you cannot just plug into the difference quotient (which is what this is called) ever

OpenStudy (misty1212):

if you can use all your calculus, recognize this as the derivative of cosine, evaluated at \(\frac{3\pi}{2}\) and you do it in your head

OpenStudy (misty1212):

what is the derivative of cosine?

OpenStudy (misty1212):

@Fanduekisses did i lose you there? i hope not

OpenStudy (fanduekisses):

ok -sine(3pi/2)

OpenStudy (misty1212):

yup

OpenStudy (misty1212):

don't forget \[f'(a)=\lim_{h\to 0}\frac{f(a+h)-f(a)}{h}\] which is exactly what you have, with \[a=\frac{\pi}{2}\] and sine

OpenStudy (misty1212):

oops i mean \[a=\frac{3\pi}{2}\] but you knew that

OpenStudy (misty1212):

i guess it is not quite clear you have given to you the definition of the derivative at \(\frac{3\pi}{2}\)

OpenStudy (misty1212):

you could compute that limit by hand if you like, but you don't have to since you already know the derivative of cosine is - sine

OpenStudy (misty1212):

making your answer \[-\sin(\frac{3\pi}{2})=1\]

OpenStudy (fanduekisses):

thanks, I knew how to solve this at the beginning of the school year, idk what is going on in my head right now :(

OpenStudy (misty1212):

i bet this is one of those reviews right?

OpenStudy (fanduekisses):

yes

OpenStudy (misty1212):

just to see if you remember that stuff

OpenStudy (fanduekisses):

OOOOOOHHHHHHHHH I just remembered everything lol, or at least comprehend it =D

OpenStudy (misty1212):

whew!!!!